OpenStudy (anonymous):

Stuck on a problem factoring trinomials. The equation is a^2-4az-6z^2

OpenStudy (anonymous):

I love trinomials

OpenStudy (anonymous):

I am stuck

OpenStudy (dumbcow):

well look for factors of -6 that add up to -4 -6 +1 = -5 NO 6 -1 = 5 NO -3 + 2 = -1 NO 3 - 2 = 1 NO This is not factorable

OpenStudy (llort):

it must add up to -4 and multiply to -6

OpenStudy (anonymous):

There could be a coefficient to "a" that cancels itself out, though. Trying to figure it out...

OpenStudy (anonymous):

I know they cant be factor but it has to look something like (a-3)(a+2) something in that nature.

OpenStudy (dumbcow):

lol thats called factoring, if its prime then it can't look like that

OpenStudy (anonymous):

lol I told you I am confused hahaha.

OpenStudy (llort):

who says that it has to be a whole number... \[x = -2-\sqrt(10), y = \sqrt(10)-2\] OR \[x = \sqrt(10)-2, y = -2-\sqrt(10)\]

OpenStudy (anonymous):

You can always use the quadratic formula and solve for a. I don't see any elegant solution using whole numbers.

OpenStudy (dumbcow):

true, but usually integers are implied when factoring

OpenStudy (anonymous):

Wolfram Alpha couldn't find an answer and neither could I.

OpenStudy (anonymous):

I am sorry I wrote the equation wrong on my paper it is a^2-4az-60z^2

OpenStudy (dumbcow):

the goal isnt to solve i believe, just put into factored form :|

OpenStudy (llort):

(a-10 z) (a+6 z)

OpenStudy (llort):

it would help toget your numbers right...

OpenStudy (dumbcow):

Llort got it

OpenStudy (llort):

I always do... :P

OpenStudy (anonymous):

Sorry Llort

OpenStudy (llort):

it's ok